OpenStudy (anonymous):

The sum of the interior angles of a quadrilateral equals 360°. True False

OpenStudy (mathstudent55):

The sum of the measures of the interior angles of a polygon with n sides is (n - 2)180. Use the statement above to find whether the given statement is true or false.

OpenStudy (anonymous):

false?

OpenStudy (mathstudent55):

Let's solve this together. Look at the formula I gave you above. If a polygon has n sides, the sum of the measures of the interior is (n - 2)180. In your case, you are dealing with a quadrilateral. How many sides does a quadrilateral have?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so its true!

OpenStudy (mathstudent55):

Correct. You have n = 4. That means you get: (n - 2)180 = (4 - 2)180 = 2*180 = 360 The statement is true.